OpenStudy (anonymous):

A bag contains 7 nickels, 4 dimes, & 8 quarters. If you draw 3 coins at random from the bag, without replacement, what is the probability that you will get a nickel, a quarter, and a nickel, in that order?

OpenStudy (anonymous):

there are 19 coins total so first to get a nickel 7/19 than there are only 18 coins left so to get a quarter is 8/18 than there is only 17 left but only 6 nickels so that is 6/17 As all the events are independent the total probability is there product so 7*8*6/(19*18*17)

OpenStudy (anonymous):

my options are : 22/57, 392/6859, 7/18, 56/969, or 196/2907

OpenStudy (anonymous):

7*8*6/19*18*17=7*8/(19*3*17)=56/969

OpenStudy (anonymous):

is this your homework?